# Unlabeled Debiasing in Downstream Tasks via Class-wise Low Variance Regularization

### v1

@@ Abstract
Shahed Masoudian1, Markus Frohman1,2, Navid Rekabsaz1,3, Markus Schedl1,2 1Johannes Kepler University Linz, Austria 2Linz Institute of Technology, AI Lab, Austria 3Thomson Reuters Labs, Zug, Switzerland shahed.masoudian@jku.at
@@ 1 Introduction and Background
Several works focus on reducing the effect of these biases by improving model performance related to some specific fairness metric (empirical fairness) or by making the model blind to the existence of a certain attribute (representational fairness(Shen et al., 2022). For instance, Shen et al. (2022) leverage contrastive learning to improve empirical fairness. Recent works focus mostly on efficiency and user flexibility when it comes to debiasing using modular approaches such as sub-networks or adapters Houlsby et al. (2019). Hauzenberger et al. (2023) introduce a modular debiasing scheme with adversarial training Elazar and Goldberg (2018) and mutual information reduction Colombo et al. (2021) to control the bias in encoder LMs. Kumar et al. (2023b) use adversarial training with adapters Pfeiffer et al. (2021) to improve representational fairness. Finally, Masoudian et al. (2024) used gated adapters to improve representational fairness while preserving task performance for classification and retrieval tasks.
@@
We demonstrate the effectiveness of our method using adapters Pfeiffer et al. (2021) and two commonly used encoder LMs, BERT-Base and RoBERTa-Base. Furthermore, we show that our method, when compared to existing supervised debiasing methods, can enhance attribute removal while still showing competitive classification task performance.
@@ 2 Problem Formulation
In our initial study, we assess how much gender information can be extracted from commonly used encoder LMs. We use adapters (Pfeiffer et al., 2021, Pfeiffer variation) in combination with (Devlin et al., 2019, BERT-Base) and, additionally, a gender-debiased version of the same model (Wu et al., 2022, BERT-Base-Debias-NLI) debiased for empirical fairness, on two downstream classification tasks. We then train probes on the embeddings of the fine-tuned models to check how much information about gender can be extracted from both model variations and report the average balanced accuracy as indicator of gender information in the embeddings.
Table 1 shows the result of both models for occupation prediction (BIOS (De-Arteaga et al., 2019)) and mention prediction (PAN16 (Rangel et al., 2016)) datasets. We observed that task performance of the debiased LM, BERT-Base-Debias-NLI, is consistently lower than BERT-Base, which aligns with observations by Ghanbarzadeh et al. (2023). Moreover, training adapters contain the gender information to a great extent on BIOS; while on PAN16, BERT-Base-Debias-NLI leaks more gender information in the embeddings, although it has already been subject to debiasing.
Table 1: Result of adapter training of BERT (BERT-Base) and a gender-debiased version of BERT (AdpNLI) on two datasets with gender as the protected attribute. Here, Task corresponds to accuracy on the main classification task, and Gender is the balanced accuracy of the model concerning the protected attribute.
@@ 3 Low Variance Regularization (LVR)
$\mathcal{L}_{total}=\mathcal{L}_{t}+\lambda\mathcal{L}_{r}+\mathcal{L}_{c},$ (3)

### v3

@@ Abstract
Shahed Masoudian1, Markus Frohmann1,2, Navid Rekabsaz3, Markus Schedl1,2 1 Johannes Kepler University Linz, Austria 2 Linz Institute of Technology, AI Lab, Austria 3 Thomson Reuters Labs, Zug, Switzerland shahed.masoudian@jku.at
@@ 1 Introduction and Background
Several works focus on reducing the effect of these biases by improving model performance related to some specific fairness metric (empirical fairness) or by making the model blind to the existence of a certain attribute (representational fairness(Shen et al., 2022). For instance, Shen et al. (2022) leverage contrastive learning to improve empirical fairness. Recent works focus mostly on efficiency and user flexibility when it comes to debiasing using modular approaches such as sub-networks or adapters Houlsby et al. (2019). Hauzenberger et al. (2023) introduce a modular debiasing scheme with adversarial training Elazar and Goldberg (2018) and mutual information reduction Colombo et al. (2021) to control the bias in encoder LMs. Kumar et al. (2023b) use adversarial training with adapters Pfeiffer et al. (2021) to improve representational fairness on document classification. Finally, Masoudian et al. (2024) used gated adapters to improve representational fairness while preserving task performance for classification and retrieval tasks.
@@
We demonstrate the effectiveness of our method on document classification taskss using adapters Pfeiffer et al. (2021) and two commonly used encoder LMs, BERT-Base and RoBERTa-Base. Furthermore, we show that our method, when compared to existing supervised debiasing methods, can enhance attribute removal while still showing competitive classification task performance.
@@ 2 Problem Formulation
In our initial study, we assess how much gender information can be extracted from commonly used encoder LMs. We use adapters (Pfeiffer et al., 2021, Pfeiffer variation) in combination with (Devlin et al., 2019, BERT-Base) and, additionally, a gender-debiased version of the same model (Wu et al., 2022, BERT-NLI) debiased for empirical fairness, on two downstream classification tasks. We then train probes on the embeddings of the fine-tuned models to check how much information about gender can be extracted from both model variations and report the average balanced accuracy as indicator of gender information in the embeddings.
Table 1 shows the result of both models for occupation prediction (BIOS (De-Arteaga et al., 2019)) and mention prediction (PAN16 (Rangel et al., 2016)) datasets. We observed that task performance of the debiased LM, BERT-NLI, is consistently lower than BERT-Base, which aligns with observations by Ghanbarzadeh et al. (2023). Moreover, training adapters contain the gender information to a great extent on BIOS; while on PAN16, BERT-NLI leaks more gender information in the embeddings, although it has already been subject to debiasing.
Table 1: Result of adapter training of BERT (BERT-Base) and a gender-debiased version of BERT (BERT-NLI) on two datasets with gender as the protected attribute. Here, Task corresponds to accuracy on the main classification task, and Gender is the balanced accuracy of the model concerning the protected attribute.
@@ 3 Low Variance Regularization (LVR)
$\mathcal{L}_{total}=\mathcal{L}_{t}+\lambda\mathcal{L}_{r}+\mathcal{L}_{c}$ (3)